Please show your work. Thank you

Costs per Equivalent Unit

The following information concerns production in the Baking Department for March. All direct materials are placed in process at the beginning of production.

ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Mar. 1 Bal., 4,500 units, 4/5 completed 7,830
31 Direct materials, 81,000 units 105,300 113,130
31 Direct labor 31,200 144,330
31 Factory overhead 17,544 161,874
31 Goods finished, 82,200 units 156,000 5,874
31 Bal. ? units, 4/5 completed 5,874

a. Based on the above data, determine each cost listed below. Round "cost per equivalent unit" answers to the nearest cent.

1. Direct materials cost per equivalent unit. $
2. Conversion cost per equivalent unit. $
3. Cost of the beginning work in process completed during March. $
4. Cost of units started and completed during March. $
5. Cost of the ending work in process. $

b. Assuming that the direct materials cost is the same for February and March, did the conversion cost per equivalent unit increase, decrease, or remain the same in March?
Answer #1
Equivalent Units
Whole Units Direct Materials Conversion
Inventory in process, March 1 4500 0 900 =4500*1/5
Started and completed in March 77700 77700 77700 =82200-4500
Transferred to Finished goods in March 82200 77700 78600
Inventory in process, March 31 3300 3300 2640 =3300*4/5
Total units to be assigned costs 85500 81000 81240
Cost Information
Costs per equivalent unit:
Direct Materials Conversion
Total costs for March in Baking Department 105300 48744 =31200+17544
Total equivalent units 81000 81240
Cost per equivalent unit 1.30 0.60
Costs charged to production:
Direct Materials Conversion Total
Inventory in process, March 1 7830
Costs incurred in March 154044
Total costs accounted for by the Baking Department 161874
Cost allocated to completed and partially completed units:
Inventory in process, March 1 balance 7830
To complete inventory in process, March 1 0 540 540
Cost of completed March 1 work in process 8370
Started and completed in March 101010 46620 147630
Transferred to Finished goods in March 156000
Inventory in process, March 31 4290 1584 5874
Total costs assigned by the Baking Department 161874
a
1. Direct materials cost per equivalent unit $1.30
2. Conversion cost per equivalent unit $0.60
3. Cost of beginning work in process completed during March $8370
4. Cost of units started and completed during March $147630
5. Cost of the ending work in process $5874
b
Conversion cost per equivalent unit prior period = (7830-4500*1.30)/(4500*4/5)= $0.55
Conversion cost per equivalent unit increase